In a class consisting of 5 females and 10 males, two students are selected to present a homework paper. What is the probability that both will be female?
Answer by stanbon(75887)

In a class consisting of 5 females and 10 males, two students are selected to present a homework paper. What is the probability that both will be female?
-----
# of ways to select 2 females: 5C2 = 10
---
# of random pairs:: 15C2 = (15*14)/(1*2) = 105
----
P(pick 2 females) = 10/105 = 2/21
-----------------
Cheers,
Stan H.
=================